OpenStudy (anonymous):

How is y=3sec(2x) related to y=sec(x)?

OpenStudy (anonymous):

Like what is the horizontal/vertical compression/shifts and by how many units?

OpenStudy (zpupster):

The formula which is used to graph the sec function: y=Asec(Bx-C), Amplitude=A, period=2π/B, Phase-shift=C/B.

OpenStudy (anonymous):

Thank you!